OpenStudy (anonymous):

Please Help Multiply, if possible. Then simplify. square root sign -5 times the square root sign 5

OpenStudy (anonymous):

\[\sqrt{5}\times \sqrt{-5}\]

OpenStudy (precal):

\[\sqrt{-25}\] \[\pm5i\]